OpenStudy (anonymous):

Find the interval of continuity and differentiability of the following function: f(x) = (\sqrt (x^2 - x)) + (\frac (1 / ln(2x - 1))) Step by step guide would be best

OpenStudy (anonymous):

\[f(x)=\sqrt{x^2-x}+\frac{1}{\ln (2x-1)}\]

OpenStudy (anonymous):

yh thats the function

OpenStudy (anonymous):

under radical must be greater than or equal to zero so firstly we must have : \[x^2-x=x(x-1)\ge0\]

OpenStudy (anonymous):

and for the fraction?

OpenStudy (anonymous):

consider the ristrictions for input of \(\ln\) function : \[2x-1>0\]and dont forget to figure out when division by zero takes place\[\ln(2x-1)\neq0\]\[2x-1\neq1\]

OpenStudy (anonymous):

so what are you saying the interval of continuity is?

OpenStudy (anonymous):

i would say \[(1,\infty)\]

OpenStudy (anonymous):

and with regards to differentiability?